OpenStudy (anonymous):

determine the exact value of sec 3pi/2

OpenStudy (jdoe0001):

\(\bf sec(\theta)\implies \cfrac{1}{cos(\theta)}\quad thus\quad sec\left(\frac{3\pi}{2}\right)\implies \cfrac{1}{cos\left(\frac{3\pi}{2}\right)}\)